1. A stock has an expected return of 9 percent, its beta is 0.5, and the...

1.

A stock has an expected return of 9 percent, its beta is 0.5, and the risk-free rate is 3.6 percent. What must the expected return on the market be?


rev: 09_20_2012

10.80%

15.12%

13.68%

14.98%

14.40%

2.

You own the following portfolio of stocks. What is the portfolio weight of stock C?

Stock Number of Shares Price per Share
A 120 $38
B 810 $34
C 450 $58
D 260 $57

35.74 percent

20.30 percent

37.72 percent

64.26 percent

58.02 percent
Answer #1

1.Expected return=risk-free rate +Beta*(market rate- risk-free rate )

9=3.6+0.5*(Market rate-3.6)

(9-3.6)/0.5=Market rate-3.6

Market rate=(9-3.6)/0.5+3.6

=14.4%

2.

Total value=(Number of shares*Price per share)
A (120*38)=$4560
B (810*34)=$27540
C (450*58)=$26100
D (260*57)=$14820
Total=$73020

Hence portfolio weight of C=(26100/73020)

=35.74%(Approx).